OpenStudy (anonymous):

Find the derivative of with respect to x: Y=-2x^3 sin(4x-1)

OpenStudy (anonymous):

if y= a(x)*b(x) dy/dx= a(x)*b'(x)+a'(x)*b(x)

OpenStudy (amistre64):

u= 4x-1 d(sin(u)) = cos(u) d(4x-1) = 4 Dx(sin(4x-1)) = 4 cos(4x-1) you can use that in your product rule as b'(x) if it helps....